Find the slope of the tangent line at the indicated point
Ycubed plus y is equal to xsquared at (1,3)
Write an equation of the tangent line.
1
Expert's answer
2020-06-03T18:35:53-0400

Equation of the Curve :

"y^{3} + y = x^{2}"


To find equation of tangent, first we need to find slope of tangent at (1, 3) i.e. "\frac{dy}{dx}"

So, Differentiating both sides with respect to x,

"3y^{2}{\frac{dy}{dx}} + {\frac{dy}{dx}} = 2x"


"(3y^{2} + 1){\frac{dy}{dx}} = 2x"


"{\frac{dy}{dx}} = {\frac{2x}{3y^{2}+1}}"


At (1,3) slope, m = "{\frac{dy}{dx}} = {\frac{2}{28}} = {\frac{1}{14}}"


Hence, equation of the tangent is given by


"(y - 3) = {\frac{1}{14}}(x - 1)"


"y = {\frac{1}{14}}(x + 41)"